OpenStudy (katherinesmith):

simplify:

OpenStudy (katherinesmith):

\[(\frac{ 5x ^{-1} }{ x ^{3} })^{^{-1}}\]

OpenStudy (katherinesmith):

@Hero

hero (hero):

First understand that negative exponent means "inverse". So if you have for example \(\large 5^{-1}\) you read it as the inverse of \(\large 5\). The inverse of \(\large5\) is \(\large\frac{1}{5}\) so \(\large5^{-1} = \frac{1}{5}\) And in general: \[x^{-a} = \frac{1}{x^a}\] In other words, the inverse of x to the power -a, is one over x to the power a.

OpenStudy (katherinesmith):

oh my god. that's confusing.

OpenStudy (katherinesmith):

x^3/5 ?

hero (hero):

I can show you in vyew.

OpenStudy (katherinesmith):

okay, i don't have an account but i can make one

OpenStudy (katherinesmith):

my flash player isn't up to date so it's not gonna work

OpenStudy (katherinesmith):

haha don't worry about it i'll just move on to another question. thanks anyways

hero (hero):

You don't have to make an account. I already have one.